A simple random sample of 60 items resulted in a sample mean of 73. The population standard deviation is 15.

a. Compute the 95% confidence interval for the population mean (to 1 decimal).

(____) , (____)

b. Assume that the same sample mean was obtained from a sample of 120 items. Provide a 95% confidence interval for the population mean (to 2 decimals).

(___) , (___)

c. What is the effect of a larger sample size on the margin of error?
Select- (increases), (It decreases), (It stays the same), (it cannot be determined from the given data)

I NEED IT ASAP, thanks
Answer #1

a)

For n = 60,

95% confidence interval for \mu is

\bar{x} - Z\alpha/2 * \sigma / sqrt(n) < \mu < \bar{x} + Z\alpha/2 * \sigma / sqrt(n)

73 - 1.96 * 15 / sqrt(60) < \mu < 73 + 1.96 * 15 / sqrt(60)

69.2 < \mu < 76.8

95% CI is (69.2 , 76.8)

b)

For n = 120,

95% confidence interval for \mu is

\bar{x} - Z\alpha/2 * \sigma / sqrt(n) < \mu < \bar{x} + Z\alpha/2 * \sigma / sqrt(n)

73 - 1.96 * 15 / sqrt(120) < \mu < 73 + 1.96 * 15 / sqrt(120)

70.32 < \mu < 75.68

95% CI is (70.32 , 75.68)

c)

Larger sample size decreses the margin of error